OpenStudy (anonymous):

HELP MEDAL! Using a directrix of y = -3 and a focus of (2, 1), what quadratic function is created?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@freckles

OpenStudy (freckles):

I like deriving the equation for the parabola each time The distance between the point (x,y) to (2,1) should have the same distance between (x,y) to (x,-3) This is because the definition of a parabola says for all points (x,y) on the parabola we should have the distance between (x,y) to the focus is the same as (x,y) to the directrix

OpenStudy (freckles):

Use the distance formula to find the distance between (x,y) and (2,1) and use the distance formula to find the distance between (x,y) and (x,-3)

OpenStudy (anonymous):

so how do you do that

OpenStudy (tkhunny):

directrix of y = -3 focus of (2, 1) Immediately, you should see: vertex: (2,-1) "p": 2 Now, just write down the equation.

OpenStudy (anonymous):

how i dont get this part?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Vertex form of a parabola is what? Write it out.

OpenStudy (anonymous):

y=a(x-h)^2+k

OpenStudy (tkhunny):

Okay, fine. That is the Vertex Form, but that is not the way to go about it. Not enough information about "a". The "Conic" form is this. 4p(y-k) = (x-h)^2 h and k are exactly the same as in the Vertex From. p is the distance from Vertex to Focus or the same distance from Vertex to Directrix. Note: This new "4p" from the conic form is the same as 1/a from the Vertex Form.

OpenStudy (anonymous):

f(x) = one eighth (x - 2)2 - 1 f(x) = -one eighth (x + 2)2 - 1 f(x) = one half (x - 2)2 + 1 f(x) = -one half (x + 2)2 - 1 any of these

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@freckles got any help

OpenStudy (freckles):

do you want to derive the equation with the distance formula or just use what tk said

OpenStudy (freckles):

the p should equal the distance between the vertex and directrix and the sign of p should be determined by if the parabola is faced up or down

OpenStudy (freckles):

 vertex falls halfway between the directrix and the focus